X(0) -90020 4) Solve for lups - [-3x K te 42 12x7 at
Answer #1

dx dt 1-3 X dx 1-3% = dt integrate it we get In 21-3x1 = tt (ज) cult X (6) có at t=0, XCO so fea So eq. (1) becomes = 1/2 Init is given that dy - 2+ 2x -Y dy + Y 2 + 2x dt MT dy at + y = 2+ ? (1-etty (: eq@ 20 T dy at +y 2 -3t 63) [1dt = e (first oret y Lt) & et 2 -2E e E2) tc 3 zt + + ē + c (4) u sy . 100 so & + te 3 W w 8 e so es (4) becomes et ya) . & et + + 6 24 - 3 &